# Supervised heterogeneous network estimation via survival-based Bayesian graphical models

### v2

@@ Abstract
Heterogeneity is a fundamental characteristic of cancer. To accommodate heterogeneity, subgroup identification has been extensively studied and broadly categorized into unsupervised and supervised analysis. Compared to unsupervised analysis, supervised approaches potentially hold greater clinical implications. Under the unsupervised analysis framework, several methods focusing on network-based subgroup identification have been developed, offering more comprehensive insights than those restricted to mean, variance, and other simplistic distributions by incorporating the interconnections among variables. However, research on supervised network-based subgroup identification remains limited. In this study, we develop a novel supervised Bayesian graphical model for jointly identifying multiple heterogeneous networks and subgroups. In the proposed model, heterogeneity is not only reflected in molecular data but also associated with a clinical outcome, and a novel similarity prior is introduced to effectively accommodate similarities among the networks of different subgroups, significantly facilitating clinically meaningful biological network construction and subgroup identification. The consistency properties of the estimates are rigorously established, and an efficient algorithm is developed. Extensive simulation studies and a real-world application to TCGA data are conducted, which demonstrate the advantages of the proposed approach in terms of both subgroup and network identification.
@@ 1 Introduction
For cancer, heterogeneity is a fundamental characteristic and is essential for unraveling the inherent complexities of biological systems. To describe and accommodate heterogeneity, researchers have conducted extensive studies on subgroup identification with high-dimensional molecular data, which can be roughly classified into unsupervised and supervised. Unsupervised approaches typically emphasize differences in the means of molecular factors (Table 1, a), with examples including [18] and [4], and variances (Table 1, b), as exemplified by [26]. Under the supervised analysis framework, a popular approach is the finite mixture of regression models, which captures the heterogeneity of genetic effects on clinical outcomes (Table 1, d). Examples include both Bayesian approaches [21, 1] and penalization approaches [24, 31]. These models enable the integration of information from both clinical outcomes and molecular data to identify clinically relevant subgroups. Despite significant achievements, existing techniques remain inadequate as they focus on the distributional properties or the influences of individual variables but neglect the interconnections among variables. There is a strong need for more informative approaches for subgroup identification.
Table 1: Comparison of unsupervised and supervised subgroup identification models.
@@
Unsupervised a. Mean-based: $\bm{X}\sim\mathcal{N}(\bm{\mu}_{k},\bm{V})$, where $\bm{\mu}_{k}$ is the mean vector in the $k$th subgroup and $\bm{V}$ is the diagonal covariance matrix. Example: $\bm{\mu}_{1}=(1,0,0)^{\top},\ \bm{\mu}_{2}=(0,1,-1)^{\top}$.
@@
b. Variance-based: $\bm{X}\sim\mathcal{N}(\bm{\mu}_{k},\bm{D}_{k})$, where $\bm{D}_{k}$ is the diagonal covariance matrix in the $k$th subgroup. Example: $\bm{D}_{1}=\text{diag}(1,2,1),\ \bm{D}_{2}=\text{diag}(2,1,2)$.
@@
c. Network-based: $\bm{X}\sim\mathcal{N}(\bm{\mu}_{k},\bm{\Omega}_{k}^{-1})$, where $\bm{\Omega}_{k}$ is the precision matrix in the $k$th subgroup. Example: $\bm{\Omega}_{1}=\begin{pmatrix}1&-1&0\\ -1&2&0\\ 0&0&1\end{pmatrix},\ \bm{\Omega}_{2}=\begin{pmatrix}2&0&-1\\ 0&1&0\\ -1&0&2\end{pmatrix}$.
@@
Supervised d. Linear effect-based: $Z\sim\mathcal{N}(\bm{\beta}_{k}^{\top}\bm{X},\sigma_{k}^{2})$, where $\bm{\beta}_{k}$ and $\sigma_{k}$ are the regression parameter vector and standard deviation in the $k$th subgroup. Example: $\bm{\beta}_{1}=(-1,0,0)^{\top},\ \bm{\beta}_{2}=(0,1,-1)^{\top};\ \sigma_{1}=1,\ \sigma_{2}=2$.
@@
The importance of network-based subgroup identification has been emphasized in recent literature [22]. Subtype-specific networks are pivotal in understanding complex biological processes. They utilize information embedded in the interconnections among variables (Table 1, c), providing richer insights than marginal means and variances. As illustrated in the literature [7, 19], in addition to heterogeneity, there can be commonalities across subgroup-specific networks, which can help improve estimation efficiency and enhance interpretability through information borrowing.
@@
Among the diverse statistical methods for network analysis, Gaussian graphical models (GGMs) have been prominent. Here, the nodes of a network are represented by random variables with a joint Gaussian distribution, while the edges reflect conditional dependencies between nodes that are determined by non-zero entries of the inverse covariance matrix (or precision matrix) [16, 10, 5]. To detect unknown heterogeneity, several network-based methods have been proposed under the framework of the Gaussian mixture model (GMM), with particular attention to identifying both structural differences and potential commonalities among multiple networks. For example, [11] incorporated a joint fused graphical lasso penalty into GMM for learning the related structures across different unknown subgroups. Similarly, a SCAN method proposed by [12] used a joint group graphical lasso penalty for simultaneous subgroup identification and joint graphical model estimation. [20] proposed decomposing the GGM into a set of sparse regression problems and introduced a composite minimax concave penalty for jointly estimating multiple networks. Under the Bayesian framework, [6] proposed a BCGS method that imposes a joint spike-and-slab graphical lasso prior on the precision matrices.
@@
However, these GMM-based approaches belong to the unsupervised paradigm, focusing solely on the discovery of molecularly separable subgroups, do not consider any clinical outcomes, and may lead to subgroups lacking sufficient clinical implications. Published literature has demonstrated that supervised heterogeneity analysis that incorporates clinical information is promising in identifying clinically relevant subgroups and provides a more comprehensive understanding of disease [24, 31]. However, the existing supervised heterogeneity methods predominantly focus on detecting heterogeneous predictor-outcome associations without appropriately accommodating the network structures among predictors. Overall, there is a notable gap in jointly exploring molecular network structures and conducting subgroup identification in a supervised manner.
@@
In this paper, we develop a novel supervised network-based heterogeneity analysis approach to simultaneously conduct network estimation and subgroup identification. This approach is among the first to conduct joint estimation for multiple networks in a supervised manner with unknown subgroup structures. Significantly advancing from the existing unsupervised network-based subgroup identification approaches [12, 14, 6], we develop a novel Bayesian framework integrating a finite mixture of Gaussian graphical models with supervised regression models to accommodate heterogeneity involving not only the distributions of predictors but also the predictor-outcome associations, leading to clinically more meaningful biological network construction and subgroup identification. Additionally, advancing from the existing joint network analysis based on the group penalties [7, 12, 14, 6], a novel similarity prior is introduced to capture the similarity information between the values of connection effects of distinct networks, facilitating more effective information borrowing and more appropriate modeling of the underlying biological mechanisms. Furthermore, unlike the existing supervised subgroup identification approaches [1, 31], we fully take the specific genetic networks into account while simultaneously accommodating the clinical information. The consistency properties of the proposed estimators are rigorously established, achieved through an efficient algorithm. Comprehensive simulation studies are conducted, demonstrating the advantages of the proposed approach in terms of subgroup and network identification. Additionally, a real-world application to TCGA data highlights the practical utilization of our approach.
@@ 2 Methods
We consider data with a survival outcome. Modeling survival presents more challenges due to non-negativity and censoring, and similar approaches can be applied to other clinical outcomes. Let $T$, $\Delta$, and $\bm{X}=\left(X_{1},\ldots,X_{p}\right)^{\top}$ represent the logarithm of the minimum of survival time and censoring time, censoring indicator, and $p$-dimensional vector of predictors (e.g., gene expressions), respectively. Here, $T=\operatorname{min}\left(Z,C\right)$ with $Z$ and $C$ being the logarithms of survival time and censoring time, and $\Delta={I}\{Z\leq C\}$ with ${I}\{\cdot\}$ being the indicator function. Assume there are $n$ subjects $\left\{\left(t_{i},\delta_{i},\bm{x}_{i}\right),i=1,\ldots,n\right\}$, which form $K$ subgroups. Denote $\bm{z},\bm{c},\bm{\delta},$ and $\bm{t}$ as the vectors of $z_{i}$’s, $c_{i}$’s, $\delta_{i}$’s, and $t_{i}$’s, respectively, and X as the $n\times p$ matrix composed of $\bm{x}_{i}$’s.
@@ 2.1 Bayesian mixture modeling
[…]
@@
where $\ddot{\beta}_{0k}$ and $\ddot{\bm{\beta}}_{k}=\left(\ddot{\beta}_{k,1},\ldots,\ddot{\beta}_{k,p}\right)^{\top}$ are the subgroup-specific intercept and regression parameters, and $\epsilon_{i}$ is a normally distributed random variable with mean 00 and precision parameter ${\tau}_{k}^{2}$. We propose the following Bayesian mixture model with a reparameterization technique:
$t_{i}\mid\left(\delta_{i},z_{i},c_{i}\right)=z_{i}^{\delta_{i}}c_{i}^{1-\delta_{i}},\delta_{i}\mid\left(z_{i},c_{i}\right)={I}\left\{z_{i}\leq c_{i}\right\},c_{i}\sim f_{C}(c_{i}),$ (1)
$z_{i}\mid\left(\bm{\beta},\bm{x}_{i},g_{i}=k;\bm{\tau}\right)\sim\mathcal{N}\left(z_{i}\mid\beta_{0k}/\tau_{k}+\bm{\beta}_{k}^{\top}\bm{x}_{i}/\tau_{k},\tau_{k}^{-2}\right),\bm{x}_{i}\mid\left(\bm{\mu},\bm{\Omega},g_{i}=k\right)\sim\mathcal{N}\left(\bm{x}_{i}\mid\bm{\mu}_{k},\bm{\Omega}_{k}^{-1}\right).$
@@
Here, $f_{C}$ is the density function of the logarithm of censoring time $C$, $\bm{\mu}_{k}=\left(\mu_{k,1},\cdots,\mu_{k,p}\right)^{\top}$ and $\bm{\Omega}_{k}=\left(\omega_{k,jl}\right)_{p\times p}$ are the mean and precision matrix for the predictors in the $k$th subgroup, and $\bm{\beta}$, $\bm{\tau}$, $\bm{\mu}$, and $\bm{\Omega}$ are the vectors of $\left({\beta}_{0k},\bm{\beta}_{k}\right)$’s, ${\tau}_{k}$’s, $\bm{\mu}_{k}$’s, and $\bm{\Omega}_{k}$’s, respectively, with ${\bm{\beta}}_{k}=\ddot{\bm{\beta}}_{k}\tau_{k}$ and $\beta_{0k}=\ddot{\beta}_{0k}\tau_{k}$. For the $k$th subgroup, we estimate the network based on $\bm{\Omega}_{k}$, where $\omega_{k,jl}\neq 0$ corresponds to a connection between predictors $l$ and $j$.
@@
After integrating out $\bm{z}$ and $\bm{c}$, we have the joint distribution of $\left(\bm{t},\bm{\delta},\textbf{X}\right)$ as:
@@
$f\left(\bm{t},\bm{\delta},\textbf{X}\mid\bm{\beta},\bm{\mu},\bm{\Omega},\bm{\tau},\bm{g}\right)=\prod_{i=1}^{n}\prod_{k=1}^{K}\left[f_{k}\left(t_{i},\bm{x}_{i},\delta_{i}\mid{\bm{\mu}}_{k},{\bm{\Omega}}_{k},{\bm{\beta}}_{k},{\beta}_{0k},{{\tau}}_{k}\right)\right]^{I\left\{g_{i}=k\right\}}$
$=$ $\prod_{i=1}^{n}\prod_{k=1}^{K}\left\{f_{\bm{X}}\left(\bm{x}_{i}\mid\bm{\mu}_{k},\bm{\Omega}_{k}\right)\left[f_{Z}\left(t_{i}\mid\bm{x}_{i},\bm{\beta}_{k},\beta_{0k},{\tau}_{k}\right)S_{C}(t_{i})\right]^{\delta_{i}}\left[f_{C}(t_{i})S_{Z}\left(t_{i}\mid\bm{x}_{i},\bm{\beta}_{k},\beta_{0k},{\tau}_{k}\right)\right]^{1-\delta_{i}}\right\}^{{I}\left\{g_{i}=k\right\}},$
where $f_{\bm{X}}$ and $f_{Z}$ are the density functions of $\bm{X}$ and $Z$ given in (1), respectively, and $S_{C}$ and $S_{Z}$ are the survival functions of $C$ and $Z$, respectively.
@@
In (1), we assume random censoring with $c_{i}$ being independent of $z_{i}$ and $\bm{x}_{i}$, as in the literature [24]. We propose combining the AFT model with the mixture Gaussian graphical model to accommodate the multi-layer heterogeneity. Specifically, in the AFT model, we introduce a subgroup-specific coefficient $\bm{\beta}_{k}$ to explicitly capture the heterogeneous associations of the molecular predictors with survival time. Here, a reparameterization technique based on the noise level is adopted for $\bm{\beta}_{k}$ to ensure that the proposed estimator remains scale-invariant under the mixture model [25, 31]. This technique facilitates an “adaptive adjustment” of the coefficients, aligning them with data characteristics and leading to more accurate estimations. On the other hand, within the Gaussian graphical model, we utilize $\bm{\mu}_{k}$ and $\bm{\Omega}_{k}$ to characterize heterogeneity attributed to both differential mean expressions and differential network structures. Consequently, the network’s heterogeneity is not only intricately linked to the interested measurements but also closely connected to the distribution of survival time, making the resulting networks clinically more meaningful. For example, researchers can better understand the differences between patients with good survival and those with bad survival on specific biological networks, thereby more effectively guiding subsequent treatments. Here, to simplify our approach, consistent with prior studies [12, 8], we assume that the subgroup number $K$ is identifiable.

### v4

# Supervised Bayesian Joint Graphical Model for Simultaneous Network Estimation and Subgroup Identification

@@ Abstract
Heterogeneity is a fundamental characteristic of cancer. To accommodate heterogeneity, subgroup identification has been extensively studied and broadly categorized into unsupervised and supervised analysis. Compared to unsupervised analysis, supervised approaches potentially hold greater clinical implications. Under the unsupervised analysis framework, several methods focusing on network-based subgroup identification have been developed, offering more comprehensive insights than those restricted to mean, variance, and other simplistic distributions by incorporating the interconnections among variables. However, research on supervised network-based subgroup identification remains limited. In this study, we develop a novel supervised Bayesian graphical model (SBJGM) for jointly identifying multiple heterogeneous networks and subgroups. In the proposed model, heterogeneity is not only reflected in molecular data but also associated with a clinical outcome, and a novel similarity prior is introduced to effectively accommodate similarities among the networks of different subgroups, significantly facilitating clinically meaningful biological network construction and subgroup identification. The consistency properties of the estimates are rigorously established, and an efficient algorithm is developed. Extensive simulation studies and a real-world application to The Cancer Genome Atlas (TCGA) data are conducted, which demonstrate the advantages of the proposed approach in terms of both subgroup and network identification.
@@ 1 Introduction
For cancer, heterogeneity is a fundamental characteristic and is essential for unraveling the inherent complexities of biological systems. To describe and accommodate heterogeneity, researchers have conducted extensive studies on subgroup identification with high-dimensional molecular data, which can be roughly classified into “unsupervised” and “supervised”. In this paper, the terms “unsupervised” and “supervised” are defined based on the type of information used. “Unsupervised” means that the method relies solely on the distributions of molecular variables, whereas “supervised” means that the method leverages predictor-outcome associations to inform heterogeneity modeling. More specifically, unsupervised approaches typically emphasize differences in the means of molecular factors $\boldsymbol{X}$ (Table 1, a), with examples including Pan and Shen (2007) and Chang et al. (2018), and variances (Table 1, b), as exemplified by Xie et al. (2008). Under the supervised analysis framework, a popular approach is the finite mixture of regression models, which captures the heterogeneity of genetic effects on clinical outcome $Z$ (Table 1, d). Examples include both Bayesian approaches (Raman et al., 2010; Ahmad and Fröhlich, 2017) and penalization approaches (Shokoohi et al., 2019; Zhong et al., 2023). These models enable the integration of information from both clinical outcomes and molecular data to identify clinically relevant subgroups. Despite significant achievements, existing techniques remain inadequate as they focus on the distributional properties or the influences of individual variables but neglect the interconnections among variables. There is a strong need for more informative approaches for subgroup identification.
Table 1: Comparison of unsupervised and supervised subgroup identification models, where $\boldsymbol{X}$ denotes the molecular variables and $Z$ denotes the clinical outcome.
@@
Unsupervised (relying solely on the heterogeneous distributions of $\boldsymbol{X}$) a. Mean-based: $\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{V})$, where $\boldsymbol{\mu}_{k}$ is the mean vector in the $k$th subgroup and $\boldsymbol{V}$ is the diagonal covariance matrix. Example: $\boldsymbol{\mu}_{1}=(1,0,0)^{\top},\ \boldsymbol{\mu}_{2}=(0,1,-1)^{\top}$.
@@
b. Variance-based: $\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{D}_{k})$, where $\boldsymbol{D}_{k}$ is the diagonal covariance matrix in the $k$th subgroup. Example: $\boldsymbol{D}_{1}=\text{diag}(1,2,1),\ \boldsymbol{D}_{2}=\text{diag}(2,1,2)$.
@@
c. Network-based: $\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{\Omega}_{k}^{-1})$, where $\boldsymbol{\Omega}_{k}$ is the precision matrix in the $k$th subgroup. Example: $\boldsymbol{\Omega}_{1}=\begin{pmatrix}1&-1&0\\ -1&2&0\\ 0&0&1\end{pmatrix},\ \boldsymbol{\Omega}_{2}=\begin{pmatrix}2&0&-1\\ 0&1&0\\ -1&0&2\end{pmatrix}$.
@@
Supervised (leveraging heterogeneous associations between $Z$ and $\boldsymbol{X}$) d. Linear effect-based: $Z\sim\mathcal{N}(\boldsymbol{\beta}_{k}^{\top}\boldsymbol{X},\sigma_{k}^{2})$, where $\boldsymbol{\beta}_{k}$ and $\sigma_{k}$ are the regression parameter vector and standard deviation in the $k$th subgroup. Example: $\boldsymbol{\beta}_{1}=(-1,0,0)^{\top},\ \boldsymbol{\beta}_{2}=(0,1,-1)^{\top};\ \sigma_{1}=1,\ \sigma_{2}=2$.
@@
The importance of network-based subgroup identification has been emphasized in recent literature (Ren et al., 2022). Subtype-specific networks are pivotal in understanding complex biological processes. They utilize information embedded in the interconnections among variables (Table 1, c), providing richer insights than marginal means and variances. As illustrated in the literature (Danaher et al., 2014; Peterson et al., 2015), in addition to heterogeneity, there can be commonalities across subgroup-specific networks, which can help improve estimation efficiency and enhance interpretability through information borrowing.
@@
Among the diverse statistical methods for network analysis, Gaussian graphical models (GGMs) have been prominent. Here, the nodes of a network are represented by random variables with a joint Gaussian distribution, while the edges reflect conditional dependencies between nodes that are determined by non-zero entries of the inverse covariance matrix (or precision matrix) (Gan et al., 2019; Colombi et al., 2024). To detect unknown heterogeneity, several network-based methods have been proposed under the framework of the Gaussian mixture model (GMM), with particular attention to identifying both structural differences and potential commonalities among multiple networks. For example, Gao et al. (2016) incorporated a joint fused graphical lasso penalty into GMM for learning the related structures across different unknown subgroups. Similarly, a Simultaneous Clustering And estimatioN (SCAN) method proposed by Hao et al. (2018) used a joint group graphical lasso penalty for simultaneous subgroup identification and joint graphical model estimation. Qin et al. (2024) proposed decomposing the GGM into a set of sparse regression problems and introduced a composite minimax concave penalty for jointly estimating multiple networks. Under the Bayesian framework, Dai and Jin (2022) proposed a Bayesian Clustering and multiple Graphical Selection (BCGS) method that imposes a joint spike-and-slab graphical lasso prior on the precision matrices.
@@
However, these GMM-based approaches belong to the unsupervised paradigm, focusing solely on the discovery of molecularly separable subgroups, do not consider any clinical outcomes, and may lead to subgroups lacking sufficient clinical implications. Published literature has demonstrated that supervised heterogeneity analysis that incorporates clinical information is promising in identifying clinically relevant subgroups and provides a more comprehensive understanding of disease (Shokoohi et al., 2019; Zhong et al., 2023). However, the existing supervised heterogeneity methods predominantly focus on detecting heterogeneous predictor-outcome associations without appropriately accommodating the network structures among predictors. Overall, there is a notable gap in jointly exploring molecular network structures and conducting subgroup identification in a supervised manner.
@@
In this paper, we develop a novel supervised network-based heterogeneity analysis approach to simultaneously conduct network estimation and subgroup identification. This approach is among the first to conduct joint estimation for multiple networks in a supervised manner with unknown subgroup structures. Significantly advancing from the existing unsupervised network-based subgroup identification approaches (Hao et al., 2018; Li et al., 2019; Dai and Jin, 2022), we develop a novel Bayesian framework integrating a finite mixture of Gaussian graphical models with supervised regression models to accommodate heterogeneity involving not only the distributions of predictors but also the predictor-outcome associations, leading to clinically more meaningful biological network construction and subgroup identification. Additionally, advancing from the existing joint network analysis based on the group penalties (Danaher et al., 2014; Hao et al., 2018; Li et al., 2019; Dai and Jin, 2022), a novel similarity prior is introduced to capture the similarity information between the values of connection effects of distinct networks, facilitating more effective information borrowing and more appropriate modeling of the underlying biological mechanisms. Furthermore, unlike the existing supervised subgroup identification approaches (Ahmad and Fröhlich, 2017; Zhong et al., 2023), we fully take the specific genetic networks into account while simultaneously accommodating the clinical information. The consistency properties of the proposed estimators are rigorously established, achieved through an efficient algorithm. Comprehensive simulation studies are conducted, demonstrating the advantages of the proposed approach in terms of subgroup and network identification. Additionally, a real-world application to TCGA data highlights the practical utilization of our approach.
@@ 2 Methods
We consider data with a survival outcome. Modeling survival presents more challenges due to non-negativity and censoring, and similar approaches can be applied to other clinical outcomes. Let $T$, $\Delta$, and $\boldsymbol{X}=\left(X_{1},\ldots,X_{p}\right)^{\top}$ represent the logarithm of the minimum of survival time and censoring time, censoring indicator, and $p$-dimensional vector of predictors (e.g., gene expressions), respectively. Here, $T=\operatorname{min}\left(Z,C\right)$ with $Z$ and $C$ being the logarithms of survival time and censoring time, and $\Delta={I}\{Z\leq C\}$ with ${I}\{\cdot\}$ being the indicator function. Assume there are $n$ subjects $\left\{\left(t_{i},\delta_{i},\boldsymbol{x}_{i}\right),i=1,\ldots,n\right\}$, which form $K$ subgroups. Denote $\boldsymbol{z},\boldsymbol{c},\boldsymbol{\delta},$ and $\boldsymbol{t}$ as the vectors of $z_{i}$’s, $c_{i}$’s, $\delta_{i}$’s, and $t_{i}$’s, respectively, and X as the $n\times p$ matrix composed of $\boldsymbol{x}_{i}$’s.
@@ 2.1 Bayesian mixture modeling
$z_{i}=\ddot{\beta}_{0k}+\ddot{\boldsymbol{\beta}}_{k}^{\top}\boldsymbol{x}_{i}+\epsilon_{i},\text{~when~}g_{i}=k,$
@@
where $\ddot{\beta}_{0k}$ and $\ddot{\boldsymbol{\beta}}_{k}=\left(\ddot{\beta}_{k,1},\ldots,\ddot{\beta}_{k,p}\right)^{\top}$ are the subgroup-specific intercept and regression parameters, and $\epsilon_{i}$ is a normally distributed random variable with mean 0 and precision parameter ${\tau}_{k}^{2}$. We propose the following Bayesian mixture model:
$t_{i}\mid\left(\delta_{i},z_{i},c_{i}\right)=z_{i}^{\delta_{i}}c_{i}^{1-\delta_{i}},$ (1)
$\delta_{i}\mid\left(z_{i},c_{i}\right)={I}\left\{z_{i}\leq c_{i}\right\},$
$c_{i}\sim f_{C}(c_{i}),$
$z_{i}\mid\left(\boldsymbol{\beta},\boldsymbol{x}_{i},g_{i}=k;\boldsymbol{\tau}\right)\sim\mathcal{N}\left(z_{i}\mid\beta_{0k}/\tau_{k}+\boldsymbol{\beta}_{k}^{\top}\boldsymbol{x}_{i}/\tau_{k},\tau_{k}^{-2}\right),$
$\boldsymbol{x}_{i}\mid\left(\boldsymbol{\mu},\boldsymbol{\Omega},g_{i}=k\right)\sim\mathcal{N}\left(\boldsymbol{x}_{i}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Omega}_{k}^{-1}\right).$
@@
Here, $f_{C}$ is the density function of the logarithm of censoring time $C$, $\boldsymbol{\mu}_{k}=\left(\mu_{k,1},\cdots,\mu_{k,p}\right)^{\top}$ and $\boldsymbol{\Omega}_{k}=\left(\omega_{k,jl}\right)_{p\times p}$ are the mean and precision matrix for the predictors in the $k$th subgroup, and $\boldsymbol{\beta}$, $\boldsymbol{\tau}$, $\boldsymbol{\mu}$, and $\boldsymbol{\Omega}$ are the vectors of $\left({\beta}_{0k},\boldsymbol{\beta}_{k}\right)$’s, ${\tau}_{k}$’s, $\boldsymbol{\mu}_{k}$’s, and $\boldsymbol{\Omega}_{k}$’s, respectively, with ${\boldsymbol{\beta}}_{k}=\ddot{\boldsymbol{\beta}}_{k}\tau_{k}$ and $\beta_{0k}=\ddot{\beta}_{0k}\tau_{k}$. For the $k$th subgroup, we estimate the network based on $\boldsymbol{\Omega}_{k}$, where $\omega_{k,jl}\neq 0$ corresponds to a connection between predictors $l$ and $j$.
@@
After integrating out $\boldsymbol{z}$ and $\boldsymbol{c}$, we have the joint distribution of $\left(\boldsymbol{t},\boldsymbol{\delta},\textbf{X}\right)$ as:
@@
$f\left(\boldsymbol{t},\boldsymbol{\delta},\textbf{X}\mid\boldsymbol{\beta},\boldsymbol{\mu},\boldsymbol{\Omega},\boldsymbol{\tau},\boldsymbol{g}\right)=\prod_{i=1}^{n}\prod_{k=1}^{K}\left[f_{k}\left(t_{i},\boldsymbol{x}_{i},\delta_{i}\mid{\boldsymbol{\mu}}_{k},{\boldsymbol{\Omega}}_{k},{\boldsymbol{\beta}}_{k},{\beta}_{0k},{{\tau}}_{k}\right)\right]^{I\left\{g_{i}=k\right\}}$
$=$ $\prod_{i=1}^{n}\prod_{k=1}^{K}\left\{f_{\boldsymbol{X}}\left(\boldsymbol{x}_{i}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Omega}_{k}\right)\left[f_{Z}\left(t_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\beta}_{k},\beta_{0k},{\tau}_{k}\right)S_{C}(t_{i})\right]^{\delta_{i}}\left[f_{C}(t_{i})S_{Z}\left(t_{i}\mid\boldsymbol{x}_{i},\boldsymbol{\beta}_{k},\beta_{0k},{\tau}_{k}\right)\right]^{1-\delta_{i}}\right\}^{{I}\left\{g_{i}=k\right\}},$
where $f_{\boldsymbol{X}}$ and $f_{Z}$ are the density functions of $\boldsymbol{X}$ and $Z$ given in (1), respectively, and $S_{C}$ and $S_{Z}$ are the survival functions of $C$ and $Z$, respectively.
@@
In (1), we assume random censoring with $c_{i}$ being independent of $z_{i}$ and $\boldsymbol{x}_{i}$, as in the literature (Shokoohi et al., 2019). We propose combining the AFT model with the mixture Gaussian graphical model to accommodate the multi-layer heterogeneity. The log-normal AFT model is adopted for simplicity, interpretability, and ease of exposition, which is common in current practice. We introduce a subgroup-specific coefficient $\boldsymbol{\beta}_{k}$ to explicitly capture the heterogeneous associations of the molecular predictors with survival time. A reparameterization technique based on the noise level is adopted for $\boldsymbol{\beta}_{k}$ to ensure that the proposed estimator remains scale-invariant under the mixture model (Städler et al., 2010; Zhong et al., 2023). This technique facilitates an “adaptive adjustment” of the coefficients, aligning them with data characteristics and leading to more accurate estimations. On the other hand, within the Gaussian graphical model, we utilize $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Omega}_{k}$ to characterize heterogeneity attributed to both differential mean expressions and differential network structures. Consequently, the network’s heterogeneity is not only intricately linked to the interested measurements but also closely connected to the distribution of survival time, making the resulting networks clinically more meaningful. For example, researchers can better understand the differences between patients with good survival and those with bad survival on specific biological networks, thereby more effectively guiding subsequent treatments. Here, consistent with prior studies (Hao et al., 2018; Dong et al., 2024), we assume that the subgroup number $K$ is known.
Our methodological framework advances beyond existing unsupervised or single-network approaches, such as Gan et al. (2019) and Hao et al. (2018). While these prior methods do not incorporate clinical outcomes and are limited to either unsupervised mixture modeling or a single graphical model, our work introduces a supervised, multi-network approach. Methodologically, we move beyond frequentist regularization or single-graph Bayesian priors by developing a novel Bayesian hierarchical model. This model integrates censored time-to-event outcomes through an AFT model and employs a structured prior to jointly estimate multiple heterogeneous networks (Sections 2.2 and 2.3). By doing so, our approach captures heterogeneity in both predictor distributions and predictor-outcome associations, offering a more comprehensive and clinically interpretable characterization.